OpenStudy (anonymous):

Simplify the expression. (√-1)/((2-4i)-(5+3i))

OpenStudy (anonymous):

Mulitiple choice a.) (-7-3i)/58 b.) (7+3i)/40 c.) (-7+3i)/58 d.) (7-3i)/40

OpenStudy (anonymous):

\[\LARGE { \sqrt{-1}\over (2-4i)-(5+3i)}=\] \[\LARGE {i\cdot \sqrt{1}\over 2-4i-5-3i}=\] \[\LARGE {i\over -3-7i}=\] \[\LARGE {i\over -3-7i}\cdot {-3+7i \over -3+7i }=\] \[\LARGE {-3i+7i^2 \over (-3)^2-(7i)^2 }=\] \[\LARGE {-3i+7\cdot (-1 ) \over 9-49\cdot (-1) }=\] can you do it now ?.. :)

OpenStudy (anonymous):